OpenStudy (lol4112):

Medal and Fan A store is having a 10%-off sale. It gives an additional 10% off if the item is still above $100 after the first discount. A jacket costs $199.99. Will you get the second discount and if so, how many dollars will you save in total?

OpenStudy (tkhunny):

Please demonstrate the 1st 10% discount.

OpenStudy (denonakavro):

Yes, Do 199-(199/100x10) Tell me what you get

OpenStudy (lol4112):

179 if im not confused @denonakavro

OpenStudy (tkhunny):

You're missing a dime.

OpenStudy (lol4112):

Oh okay I see now

OpenStudy (denonakavro):

Will be 179.1, now find the 10 percent of 179.1 and subtractit from 179.1

OpenStudy (lol4112):

17.91?

OpenStudy (tkhunny):

Alternatively, you can go directly to the discounted price. it's a 10% discount. This means that after the discount, 100% - 10% = 90% of the price will remain. 199.00 * 0.9 = 179.10 Do it again. 179.10 * 0.9 = 161.19 You don't HAVE to go through the actual discount, but it can be useful information.

OpenStudy (denonakavro):

Yes and now subtract it from 179.1

OpenStudy (lol4112):

161.19

OpenStudy (md152727):

45

OpenStudy (tkhunny):

Notice, please, that this is NOT a 20% discount. 161.19 / 199.00 = 0.81 ==> 19% discount.

OpenStudy (lol4112):

Oh okay thanks that all I need.